The sum of two number is thirty. Twice the smaller number is four less than twice the larger number. Find the larger number
Answer by mananth(16946)

let smaller number be x
bigger number be y
x+y=30 .............1
2y-2x =4.............2

multiply (1)by -2
Multiply (2) by 1
-2x-2y= -60
-2x+2y=4

Add the two equations

-4x=-56

/-4

x=14
y=16